# The Fraction of Clumpy Galaxies in JWST Surveys over $2<z<12$

### v1

@@ Abstract
High-redshift galaxies exhibit compact regions of intense star formation, known as “clumps,” which are conspicuous in the rest-frame ultraviolet. Studying them can shed light on how they form and evolve and inform theoretical models of galaxy evolution. We examine the evolution of clumpy galaxies with redshift and stellar mass over $2<z<12$ with James Webb Space Telescope (JWST) imaging from the JWST Advanced Extragalactic Survey (JADES). Off-center clumps are detected in the rest-frame near-ultraviolet (NUV) using similar techniques to those in earlier studies based on Hubble Space Telescope (HST) images. This is done for a sample of 9,121 star-forming galaxies with stellar masses $\log\left(M_{\star}/M_{\odot}\right)\geq 8$. The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, increases from $\sim 20\%$ at $z>6.5$ to $\sim 80\%$ at $z\sim 2.75$ at $\log\left(M_{\star}/M_{\odot}\right)\geq 9$. Our $f_{\rm{clumpy}}$ values are generally higher at fixed redshift and increase faster with decreasing redshift than what studies based on HST data found, which we attribute largely to the higher sensitivity of JWST. Clumps make up $15-20\%$ of the NUV luminosity of clumpy galaxies, with little dependence on redshift or stellar mass. $f_{\rm{clumpy}}$ correlates with stellar mass and there is little correlation with specific star-formation rate. Our $f_{\rm{clumpy}}$ measurements are compared with those from simulations as well as other observations. At low redshifts ($z\lesssim 5.75$) and intermediate-to-high stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$), our results suggest gas fragmentation due to violent disk instabilities to be the dominant mechanism for forming clumps. At high redshifts and low stellar masses, compression of gas during mergers appears to dominate.
@@
software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), photutils (Bradley et al., 2020; Bradley, 2023), scipy (Virtanen et al., 2020)
1 Introduction
@@
Early Hubble Space Telescope (HST) imaging of high-redshift galaxies revealed systems with patchy, clumpy, kpc-scale substructures (Cowie et al., 1995; Giavalisco et al., 1996; van den Bergh et al., 1996; Elmegreen & Elmegreen, 2005; Elmegreen et al., 2007). Later work showed that these substructures (hereafter, “clumps”) are bright in the rest-frame ultraviolet (UV), are fairly massive ($\gtrsim 10^{7}~M_{\odot}$), have young stellar ages ($\sim 100$ Myr), and make up about $10-50\%$ of the total star-formation rates (SFRs) and about $4-20\%$ of the total stellar mass of their host galaxies, on average (Elmegreen et al., 2008, 2009; Förster Schreiber et al., 2011; Guo et al., 2012, 2015, 2018; Wisnioski et al., 2012; Wuyts et al., 2012, 2013; Livermore et al., 2015; Shibuya et al., 2016; Soto et al., 2017; Zanella et al., 2019; Mehta et al., 2021; Vanzella et al., 2021; Meštrić et al., 2022; Claeyssens et al., 2025). These results suggest that clumps were significant sites of star formation in galaxies over $0.5<z<3$, during the peak of cosmic star-formation activity (Madau & Dickinson, 2014).
@@
Studies of observed and simulated clumps have proposed two main formation mechanisms. The first is violent disk instabilities (VDI), wherein gas collapses under gravitational instabilities in marginally stable gaseous disks (Noguchi, 1998, 1999; Immeli et al., 2004a, b; Bournaud et al., 2007, 2009; Agertz et al., 2009; Dekel et al., 2009a; Ceverino et al., 2010; Romeo et al., 2010; Romeo & Agertz, 2014; Inoue et al., 2016). The VDI is sustained by the high gas fractions in high-redshift galaxies (Tacconi et al., 2020, and references therein), which are fed by the accretion of cold gas from the intergalactic medium (Birnboim & Dekel, 2003; Kereš et al., 2005, 2009; Dekel & Birnboim, 2006; Bournaud & Elmegreen, 2009; Dekel et al., 2009b, a; Nelson et al., 2013). Observational support for the VDI has been found by examining integral field spectroscopic data, with which maps of gas kinematics may be measured (Förster Schreiber et al., 2011; Genzel et al., 2011, 2014; Wisnioski et al., 2011, 2012; Newman et al., 2012; Mieda et al., 2016). The second mechanism is mergers, during which gas is compressed by tidal forces and/or turbulence due to gravitational interactions between galaxies (Di Matteo et al., 2008; Renaud et al., 2015; Nakazato et al., 2024). Observational support for mergers has come from visual inspection of the morphologies of clumpy and non-clumpy galaxies (Straughn et al., 2015) and integral field spectroscopy (Puech, 2010; Menéndez-Delmestre et al., 2013).
@@
Clumps could also evolve within their host galaxies and alter their morphologies. Theoretical studies and simulations have predicted that clumps experience dynamical friction and/or interactions, which cause them to migrate to the centers of their hosts and contribute to the growth of bulges (Noguchi, 1999; Immeli et al., 2004a, b; Dekel et al., 2009a; Krumholz & Dekel, 2010; Ceverino et al., 2012; Inoue & Saitoh, 2012; Perez et al., 2013; Bournaud et al., 2014; Mandelker et al., 2014, 2017; Dekel et al., 2022). However, other simulations find that clumps are short-lived and dissipate within $\sim 100$ Myr due to stellar feedback, which drives strong outflows of gas (Murray et al. 2010; Genel et al. 2012; Hopkins et al. 2012; Moody et al. 2014; Buck et al. 2017; Oklopčić et al. 2017; Meng & Gnedin 2020; see also Fensch & Bournaud 2021). In this scenario, stars formed within the clumps could contribute to the growth of the thick disk (Noguchi 1996; Inoue & Saitoh 2014; see also Bournaud et al. 2007).
@@
The evolution of the fraction of clumpy galaxies ($f_{\rm{clumpy}}$) with redshift may also provide clues to the physical processes through which clumps form and evolve (Guo et al., 2015; Shibuya et al., 2016). Using HST imaging data, $f_{\rm{clumpy}}$ has been estimated at various redshifts by identifying compact, UV-bright regions within galaxies. The behavior of $f_{\rm{clumpy}}$ with redshift is well understood at $z\leq 3$, but unclear at higher redshifts. At $z\leq 3$, $f_{\rm{clumpy}}$ peaks at $\sim 60\%$ at $z\sim 2$ and declines at lower and higher redshifts (Elmegreen et al., 2007; Overzier et al., 2009; Puech, 2010; Guo et al., 2012; Murata et al., 2014; Guo et al., 2015; Hinojosa-Goñi et al., 2016; Shibuya et al., 2016; Adams et al., 2022; Sok et al., 2022; Martin et al., 2023; Sattari et al., 2023). These studies are largely consistent with predictions from simulations of the redshift evolution of galaxies undergoing VDI, based on simulations (Cacciato et al., 2012). Support for mergers was also found at lower redshift (Puech, 2010) and lower masses (Guo et al., 2015). At $z>3$, there is disagreement on the behavior of $f_{\rm{clumpy}}$ with redshift. One group of studies found that $f_{\rm{clumpy}}$ decreases from $\sim 40\%$ at $z\sim 3$ to $\sim 20\%$ at $z\sim 8$ (Ravindranath et al., 2006; Conselice & Arnold, 2009; Oesch et al., 2010; Shibuya et al., 2016). Another found that $f_{\rm{clumpy}}$ is roughly constant with redshift at $\sim 40\%$ from $z\sim 4$ to $z\sim 8$ (Jiang et al., 2013; Kawamata et al., 2015; Curtis-Lake et al., 2016; Bowler et al., 2017; Ribeiro et al., 2017).
@@
Studies using data from the James Webb Space Telescope (JWST) have revealed a more complex picture of the clumps. Thanks to the higher angular resolution, greater sensitivity, and longer wavelength coverage of the JWST, sub-kpc substructures in faint galaxies at high redshift can now be studied in greater detail with its Near-Infrared Camera (NIRCam). JWST imaging of gravitationally lensed galaxies over $1<z<8.5$ revealed a significant population of clumps that were not detected by the HST (Claeyssens et al., 2023; Fujimoto et al., 2024). Un-lensed galaxies at $z>4$ also reveal clumps in JWST images (e.g., Tacchella et al., 2023; Hainline et al., 2024), some of which are not found in HST observations (e.g., Bik et al., 2024; Tanaka et al., 2024). On average, 76% of the rest-frame UV light from $z\simeq 6-8$ galaxies was found to come from the clumps (Chen et al., 2023), which is much higher than the $\sim 20-30\%$ measured over $0.5<z<3$ (Guo et al., 2015). The rest-optical and rest-NIR properties of clumps have been analyzed out to higher redshifts ($z\sim 2$) than what is possible with HST (Kalita et al., 2024, 2025; Mercier et al., 2025). Massive ($\log\left(M_{\star}/M_{\odot}\right)\geq 10$) star-forming galaxies at $z<3$ were found to have higher $f_{\rm{clumpy}}$ when undergoing starburst activity (Faisst et al., 2025). Finally, about 70% of spectroscopically selected UV-bright galaxies at $z\sim 7$ contain clumps (Harikane et al., 2025), in contrast with the 102010-20% inferred from pre-JWST studies (Shibuya et al., 2016, and references therein).
@@
While recent studies show that clumpy galaxies might be more abundant than expected at high redshifts, interpretation of these results is difficult due to selection effects and occasional small sample sizes. The largest post-JWST sample at $z>6$ to date for which $f_{\rm{clumpy}}$ has been measured consists of 53 galaxies (Harikane et al., 2025). That sample was selected to be intrinsically bright with absolute UV magnitudes $23<M_{UV}<-19$ mag. Tan et al. (2024) measured $f_{\rm{clumpy}}$ over $0.25<z<5$ for a sample of 909 galaxies expected to be progenitors of Milky Way analogs. Their sample was selected with bins in stellar mass that evolve with redshift. This complicates comparisons with pre-JWST studies, which largely selected galaxies with a mass range that was constant with redshift. Claeyssens et al. (2025) examined the physical properties of 1,751 clumps in 377 galaxies over $0.7<z<5.5$ that are gravitationally lensed by the cluster Abell 2744. A measurement of $f_{\rm{clumpy}}$ as a function of redshift was not performed in that work.
@@
In this paper, we measure $f_{\rm{clumpy}}$ over $2<z<12$ with JWST data for a sample of 9,121 star-forming galaxies. The selection of the sample is performed such that results from the JWST and HST data may be directly compared. This allows us to determine whether the redshift evolution of $f_{\rm{clumpy}}$ is significantly different with JWST data compared with HST data.
@@
The paper is organized as follows. In Section 2, we describe the data used in this work. The selection criteria of our sample are given in Section 3. Section 4.1 describes the completeness of the routine used to identify clumps. The redshift evolution of $f_{\rm{clumpy}}$ is discussed in Section 5. The contribution of clumps to the rest-frame NUV luminosity is shown in Section 6. In Section 7, we examine the correlation between stellar mass and sSFR with $f_{\rm{clumpy}}$. Predictions from simulations of clumpy galaxies are compared with our observations in Section 8. We discuss the implications for clump formation mechanisms in Section 9. We summarize and conclude in Section 10.
@@
Throughout this work, cosmological parameters measured by Planck Collaboration et al. (2016) are adopted, with $H_{0}=67.7~\rm{km}~\rm{s}^{-1}~\rm{kpc}^{-1},\Omega_{\rm{m}}=0.309$, and $\Omega_{\Lambda}=0.691$. All magnitudes are in the AB photometric system (Oke & Gunn, 1983).
2 Data
2.1 Images, Source Detection, and Photometry
@@
Rieke et al. (2023b) detected sources in JADES and performed photometry in nine NIRCam and 14 ground- and space-based bandpasses. Briefly, a detection image was created by stacking the F277W, F335M, F356W, F410M, and F [^4] W images and weighting them by their inverse variances. Next, various source detection routines in the python module Photutils (Bradley, 2023) were used to detect and deblend sources and perform photometry. All images are aligned to the same reference frame and drizzled to a common pixel scale of 0.″03  per pixel. Source detection and photometry are performed on HST and JWST images convolved to the F [^4] W resolution. For more details, see Section 4 in Rieke et al. (2023b).
This work relies on photometry in HST and JWST bandpasses. Spatially integrated fluxes are measured in Kron apertures with a Kron parameter of $k=2.5$. Spatially integrated errors on the HST and JWST photometry are computed as follows. Two sources of error are considered: errors due to sky subtraction and a combination of source Poisson and instrumental errors. The former are measured in apertures of various sizes that are centered on empty regions of the sky (see Section 4.2 in Rieke et al. 2023b, also Labbé et al. 2005; Whitaker et al. 2011; and Skelton et al. 2014). For JWST/NIRCam images, the latter come from the ERR extension, and for HST images, they are measured following the procedure described in Sec. 4.1 by van der Wel et al. (2012). Total errors on the JWST photometry are computed by adding these two sources of error in quadrature. For the HST images, the JADES team have measured errors due to sky subtraction and report the average values of the weight, or inverse variance, within the Kron apertures. The error due to the weight for each galaxy is computed by multiplying the average weight value by the area of the Kron aperture in pixels. Total errors on the HST photometry are calculated by adding these two errors in quadrature. A final error is added to all fluxes for each bandpass to account for zeropoint errors and mismatches in the templates between model and observed SEDs. It is equal to 5% of the flux (see, e.g., Table 4 of Dahlen et al. 2013 and Table 11 of Skelton et al. 2014).
@@
Axis ratios, half-light radii, segmentation maps, and central coordinates of galaxies were measured from the detection images by Rieke et al. (2023b). These measurements and data products are adopted for this analysis. They were calculated by applying photutils to the detection image in each field. Half-light radii were measured in circular apertures. Central coordinates were computed by taking the light-weighted centroids along the horizontal and vertical axes of the detection images.
2.2 Photometric and Spectroscopic Redshifts

### v2

# The Fraction of Clumpy Galaxies in JADES over $2<z<9$

@@ Abstract
High-redshift galaxies exhibit compact regions of intense star formation, known as “clumps,” which are conspicuous in the rest-frame ultraviolet. Studying them can shed light on how they form and evolve and inform theoretical models of galaxy evolution. We examine the evolution of clumpy galaxies with redshift and stellar mass over 2<z<9{2<z<9} with James Webb Space Telescope (JWST) imaging from the JWST Advanced Extragalactic Survey (JADES). Off-center clumps are detected in the rest-frame near-ultraviolet (NUV) using similar techniques to those in earlier studies based on Hubble Space Telescope (HST) images. This is done for a sample of 9,121 star-forming galaxies with stellar masses $\log\left(M_{\star}/M_{\odot}\right)\geq 8$. The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, increases from $\sim 10\%$} at $z\sim 7.75$} to $\sim 70\%$} at $z\sim 2.75$ at $\log\left(M_{\star}/M_{\odot}\right)\geq 9$. Our $f_{\rm{clumpy}}$ values are generally higher at fixed redshift and increase faster with decreasing redshift than what studies based on HST data found, which we attribute largely to the higher sensitivity of JWST. $f_{\rm{clumpy}}$ correlates with stellar mass. Our $f_{\rm{clumpy}}$ measurements are compared with those from simulations as well as other observations. At low redshifts ($z\lesssim 5.75$) and intermediate-to-high stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$), our results suggest gas fragmentation due to violent disk instabilities to be the dominant mechanism for forming clumps. At high redshifts and intermediate stellar masses, compression of gas during mergers appears to dominate.
@@
software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), LMFIT (Newville et al., 2020), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), photutils (Bradley et al., 2020; Bradley, 2023), scipy (Virtanen et al., 2020)
I Introduction
@@
Early Hubble Space Telescope (HST) imaging of high-redshift galaxies revealed systems with patchy, clumpy, kpc-scale substructures (Cowie et al., 1995; Giavalisco et al., 1996; van den Bergh et al., 1996; Elmegreen and Elmegreen, 2005; Elmegreen et al., 2007). Later work showed that these substructures (hereafter, “clumps”) are bright in the rest-frame ultraviolet (UV), are fairly massive ($\gtrsim 10^{7}~M_{\odot}$), have young stellar ages ($\sim 100$ Myr), and, on average, make up about $10-50\%$ of the total star-formation rates (SFRs) and about $4-20\%$ of the total stellar mass of their host galaxies (Elmegreen et al., 2008, 2009; Förster Schreiber et al., 2011; Guo et al., 2012, 2015, 2018; Wisnioski et al., 2012; Wuyts et al., 2012, 2013; Livermore et al., 2015; Shibuya et al., 2016; Soto et al., 2017; Zanella et al., 2019; Mehta et al., 2021; Vanzella et al., 2021; Meštrić et al., 2022; Claeyssens et al., 2025; Sok et al., 2025). These results suggest that clumps were significant sites of star formation in galaxies over $0.5<z<3$, during the peak of cosmic star-formation activity (Madau and Dickinson, 2014).
@@
Studies of observed and simulated clumps have proposed two main formation mechanisms. The first is violent disk instabilities (VDI), wherein gas collapses under gravitational instabilities in marginally stable gaseous disks (Noguchi, 1998, 1999; Immeli et al., 2004a, b; Bournaud et al., 2007, 2009; Agertz et al., 2009; Dekel et al., 2009b; Ceverino et al., 2010; Romeo et al., 2010; Romeo and Agertz, 2014; Inoue et al., 2016). The VDI is sustained by the high gas fractions in high-redshift galaxies (Tacconi et al., 2020, and references therein), fed by the accretion of cold gas from the intergalactic medium (Birnboim and Dekel, 2003; Kereš et al., 2005, 2009; Dekel and Birnboim, 2006; Bournaud and Elmegreen, 2009; Dekel et al., 2009a, b; Nelson et al., 2013). Observational support for the VDI comes from integral field spectroscopic data, with which maps of gas kinematics may be measured (Förster Schreiber et al., 2011; Genzel et al., 2011, 2014; Wisnioski et al., 2011, 2012; Newman et al., 2012; Mieda et al., 2016). The second mechanism is mergers, during which gas is compressed by tidal forces and/or turbulence due to gravitational interactions between galaxies (Di Matteo et al., 2008; Renaud et al., 2015; Nakazato et al., 2024). Observational support for mergers has come from visual inspection of the morphologies of clumpy and non-clumpy galaxies (Straughn et al., 2015) and integral field spectroscopy (Puech, 2010; Menéndez-Delmestre et al., 2013).
@@
Clumps could also evolve within their host galaxies and alter their morphologies. Theoretical studies and simulations have predicted that clumps experience dynamical friction and/or interactions, which cause them to migrate to the centers of their hosts and contribute to the growth of bulges (Noguchi, 1999; Immeli et al., 2004a, b; Dekel et al., 2009b; Krumholz and Dekel, 2010; Ceverino et al., 2012; Inoue and Saitoh, 2012; Perez et al., 2013; Bournaud et al., 2014; Mandelker et al., 2014, 2017; Dekel et al., 2022). However, other simulations find that clumps are short-lived and dissipate within $\sim 100$ Myr due to stellar feedback, which drives strong outflows of gas (Murray et al. 2010; Genel et al. 2012; Hopkins et al. 2012; Moody et al. 2014; Buck et al. 2017; Oklopčić et al. 2017; Meng and Gnedin 2020; see also Fensch and Bournaud 2021). Regardless whether clumps migrate to the centers of their hosts, stars formed within the clumps could also contribute to the growth of the thick disk (Noguchi 1996; Bournaud et al. 2007; Inoue and Saitoh 2014).
@@
The evolution of the fraction of clumpy galaxies ($f_{\rm{clumpy}}$) with redshift may also provide clues to the physical processes through which clumps form and evolve (Guo et al., 2015; Shibuya et al., 2016). Using HST imaging data, $f_{\rm{clumpy}}$ has been estimated at various redshifts by identifying compact, UV-bright regions within galaxies. The behavior of $f_{\rm{clumpy}}$ with redshift is well understood at $z\leq 3$, but unclear at higher redshifts. At $z\leq 3$, $f_{\rm{clumpy}}$ peaks at $\sim 60\%$ at $z\sim 2$ and declines at lower and higher redshifts (Elmegreen et al., 2007; Overzier et al., 2009; Puech, 2010; Guo et al., 2012; Murata et al., 2014; Guo et al., 2015; Hinojosa-Goñi et al., 2016; Shibuya et al., 2016; Adams et al., 2022; Sok et al., 2022; Martin et al., 2023; Sattari et al., 2023). These studies are largely consistent with predictions from simulations of the redshift evolution of galaxies undergoing VDI (Cacciato et al., 2012). Support for mergers was also found at lower redshift (Puech, 2010) and lower masses (Guo et al., 2015). At $z>3$, there is disagreement on the behavior of $f_{\rm{clumpy}}$ with redshift, with one set of observations finding that $f_{\rm{clumpy}}$ decreases from $\sim 40\%$ at $z\sim 3$ to $\sim 20\%$ at $z\sim 8$ (Ravindranath et al., 2006; Conselice and Arnold, 2009; Oesch et al., 2010; Shibuya et al., 2016), and another finding $f_{\rm{clumpy}}$ to be roughly constant with redshift at $\sim 40\%$ from $z\sim 4$ to $z\sim 8$ (Jiang et al., 2013; Kawamata et al., 2015; Curtis-Lake et al., 2016; Bowler et al., 2017; Ribeiro et al., 2017).
@@
Studies using data from the James Webb Space Telescope (JWST) have revealed a more complex picture. Thanks to the higher angular resolution, greater sensitivity, and longer wavelength coverage of the JWST, sub-kpc substructures in faint galaxies at high redshift can now be studied in greater detail with its Near-Infrared Camera (NIRCam). JWST imaging of gravitationally lensed galaxies over $1<z<8.5$ revealed a significant population of clumps that were not detected by the HST (Claeyssens et al., 2023; Fujimoto et al., 2024). Un-lensed galaxies at $z>4$ also reveal clumps in JWST images (e.g., Tacchella et al., 2023; Hainline et al., 2024), some of which are not found in HST observations (e.g., Bik et al., 2024; Tanaka et al., 2024). On average, 76% of the rest-frame UV light from $z\simeq 6-8$ galaxies was found to come from the clumps (Chen et al., 2023), which is much higher than the $\sim 20-30\%$ measured over $0.5<z<3$ (Guo et al., 2015). Finally, about 70% of spectroscopically selected UV-bright galaxies at $z\sim 7$ contain clumps (Harikane et al., 2025), in contrast with the 102010-20% inferred from pre-JWST studies (Shibuya et al., 2016, and references therein).
@@
While recent studies show that clumpy galaxies might be more abundant than expected at high redshifts, interpretation of these results is difficult due to selection effects and occasional small sample sizes. The largest post-JWST sample at $z>6$ to date for which $f_{\rm{clumpy}}$ has been measured consists of 53 galaxies (Harikane et al., 2025). That sample was selected to be intrinsically bright with absolute UV magnitudes $23<M_{UV}<-19$ mag. Claeyssens et al. (2025) examined the physical properties of 1,751 clumps in 377 galaxies over $0.7<z<5.5$ that are gravitationally lensed by the cluster Abell 2744. A measurement of $f_{\rm{clumpy}}$ as a function of redshift was not performed in that work, due to the intrinsic incompleteness of studies on lensed galaxies. Tan et al. (2025) measured $f_{\rm{clumpy}}$ over $0.25<z<5$ for a sample of 877 galaxies expected to be progenitors of Milky Way analogs. Their sample was selected with bins in stellar mass that evolve with redshift. This complicates comparisons with pre-JWST studies, which largely selected galaxies with a mass range that was constant with redshift. Sok et al. (2025) measured $f_{\rm{clumpy}}$  and stellar masses of clumps in $\approx 4,500$ lensed galaxies with magnifications $\mu\leq 2$.
@@
In this paper, we measure $f_{\rm{clumpy}}$ over 2<z<9{2<z<9} with JWST data for a sample of 9,121 star-forming galaxies. The selection of the sample is performed such that results from the JWST and HST data may be directly compared. This allows us to determine whether the redshift evolution of $f_{\rm{clumpy}}$ is significantly different between HST and JWST data.
@@
The paper is organized as follows. In Section II, we describe the data used in this work. The selection criteria of our sample are given in Section III. Section IV describes the routine used to identify clumpy galaxies and its completeness as a function of redshift and stellar mass. The redshift evolution of $f_{\rm{clumpy}}$ is discussed in Section V. In Section VI, we examine the correlation between stellar mass with $f_{\rm{clumpy}}$. Predictions from simulations of clumpy galaxies are compared with our observations in Section VII. We discuss the implications for clump formation mechanisms in Section VIII. We summarize and conclude in Section IX.
@@
Throughout this work, cosmological parameters measured by Planck Collaboration et al. (2016) are adopted, with $H_{0}=67.7~\rm{km}~\rm{s}^{-1}~\rm{kpc}^{-1},\Omega_{\rm{m}}=0.309$, and $\Omega_{\Lambda}=0.691$. All magnitudes are in the AB photometric system (Oke and Gunn, 1983).
II Data
II.1 Images, Source Detection, and Photometry
@@
Source detection and photometry in nine NIRCam and 14 ground- and space-based bandpasses in JADES was described in Rieke et al. (2023a). Briefly, a detection image was created by convolving the F277W, F335M, F356W, F410M, and F [^4] W images to the resolution of the F [^4] W image, stacking them, and weighting them by their inverse variances. Next, various source detection routines in the python module Photutils (Bradley, 2023) were used to detect and deblend sources and perform photometry. All images are aligned to the same reference frame and drizzled to a common pixel scale of 0.″03 per pixel. Photometry was then performed on HST and JWST images convolved to the F [^4] W resolution. For more details, see Section 4 in Rieke et al. (2023a).
This work relies on photometry in HST and JWST bandpasses. Spatially integrated fluxes are measured in Kron apertures with a Kron parameter of $k=2.5$. Spatially integrated errors on the HST and JWST photometry are computed as follows. Two sources of error are considered: errors due to sky subtraction and a combination of source Poisson and instrumental errors. The former are measured in apertures of various sizes that are centered on empty regions of the sky (see Section 4.2 in Rieke et al. 2023a, also Labbé et al. 2005; Whitaker et al. 2011; and Skelton et al. 2014). For JWST/NIRCam images, the latter come from the ERR extension, and for HST images, they are measured following the procedure described in Sec. 4.1 by van der Wel et al. (2012). Total errors on the JWST photometry are computed by adding these two sources of error in quadrature. For the HST images, the JADES team have measured errors due to sky subtraction and report the average values of the weight, or inverse variance, within the Kron apertures. The error due to the weight for each galaxy is computed by multiplying the average weight value by the area of the Kron aperture in pixels. Total errors on the HST photometry are calculated by adding these two errors in quadrature. A final error is added to all fluxes for each bandpass to account for zeropoint errors and mismatches in the templates between model and observed SEDs. It is equal to 5% of the flux (see, e.g., Table 4 of Dahlen et al. 2013 and Table 11 of Skelton et al. 2014).
@@
Axis ratios, half-light radii, segmentation maps, and central coordinates of galaxies were measured from the detection images by Rieke et al. (2023a). These measurements and data products are adopted for this analysis. They were calculated by applying photutils to the detection image in each field. Half-light radii were measured in circular apertures. Central coordinates were computed by taking the light-weighted centroids along the horizontal and vertical axes of the detection images.
II.2 Photometric and Spectroscopic Redshifts